One state lottery has 1,100 prizes of $1; 120 prizes of $10; 30 prizes of $50; 5 prizes of $285; 2 prizes of $1,180; and 1 prize of $2,400. Assume that 34,000 lottery tickets are issued and sold for $1. Round to 4 decimal places for the answers

What is the lottery's expected profit per ticket?

What is the lottery's standard deviation of profit per ticket?
Answer #1

here X is random variable which denote profit for a random ticket to lettery

x f(x) xP(x) x2P(x)
0 0.032353 0.00000 0.00000
-9 0.003529 -0.03176 0.28588
-49 0.000882 -0.04324 2.11853
-284 0.000147 -0.04176 11.86118
-1179 0.000059 -0.06935 81.76712
-2399 0.000029 -0.07056 169.27062
1 0.963000 0.96300 0.96300
total 0.7063 266.2663
E(x) =μ= ΣxP(x) = 0.7063
E(x2) = Σx2P(x) = 266.2663
Var(x)=σ2 = E(x2)-(E(x))2= 265.7674
std deviation=         σ= √σ2 = 16.3024

from above:

lottery's expected profit per ticket =0.7063

lottery's standard deviation of profit per ticket =16.3024